# Rectified Flows for Fast Multiscale Fluid Flow Modeling

### v1

@@ Abstract
The statistical modeling of fluid flows is very challenging due to their multiscale dynamics and extreme sensitivity to initial conditions. While recently proposed conditional diffusion models achieve high fidelity, they typically require hundreds of stochastic sampling steps at inference. We introduce a rectified flow framework that learns a time-dependent velocity field, transporting input to output distributions along nearly straight trajectories. By casting sampling as solving an ordinary differential equation (ODE) along this straighter flow field, our method makes each integration step much more effective, using as few as eight steps versus (more than) 128 steps in standard score-based diffusion, without sacrificing predictive fidelity. Experiments on challenging multiscale flow benchmarks show that rectified flows recover the same posterior distributions as diffusion models, preserve fine-scale features that MSE-trained baselines miss, and deliver high-resolution samples in a fraction of inference time.
@@ 1 Introduction
Partial differential equations (PDEs) [4] are the backbone of modeling physical phenomena, from atmospheric dynamics to aerodynamics and magnetohydrodynamics. Currently, solutions of PDEs are simulated with high-fidelity numerical solvers such as finite difference, finite element, and spectral methods [24]. These are often prohibitively expensive for many–query tasks [25] such as uncertainty quantification, optimization, and real-time prediction. In particular, these solvers encounter a significant challenge for classes of non-linear PDEs, whose solutions exhibit extreme sensitivity to initial and boundary data, leading to chaotic behavior and complex multi-scale features. In such settings, one seeks, not a deterministic trajectory, but rather the statistical solution of the PDE, i.e., the push-forward of the underlying probability measure on inputs (initial and boundary conditions etc) by the PDE solution operator [5]. Currently, these statistical solutions are sampled with ensemble Monte Carlo method by repeatedly calling high-fidelity PDE solvers [13]. Given the computational cost of the PDE solver, ensemble methods become infeasible as ensemble sizes grow.
@@
Machine learning and AI based algorithms are increasingly being explored for the data-driven simulation of PDEs [20]. In particular, neural operators [12, 2], which directly seek to learn the PDE solution operator from data, are already widely used in a PDE context [20]. Examples of neural operators have been proposed in [14, 15, 18, 26, 11, 22, 31, 10, 16, 1] and references therein.
@@
However, these neural operators have been found to be unsuitable for simulating PDE with chaotic, multiscale solutions such as governing equations of fluid dynamics. For instance, recently in [21], it was shown both theoretically and empirically that neural operators, trained to minimize mean absolute (square) errors, will always collapse to the ensemble mean when used for simulating statistics of PDEs with chaotic multiscale solutions.
@@
Given this context, generative AI models—especially score-based and diffusion models—have recently shown remarkable success at learning complex, high-dimensional distributions by progressive denoising score matching  [29]. In the context of chaotic PDEs, GenCFD [21] introduced a conditional diffusion framework that preserves spectral content across fine scales, mitigating the “spectral collapse” of naive MSE-trained surrogates, see also [23, 7, 8, 9]. However, sampling a score-based Diffusion model requires simulating a reverse-time stochastic differential equation (SDE) over 100+ discretization steps, each invoking a large ML model on a moderate-to-high resolution spatial grid, resulting in high runtimes that undermine real-time or resource-constrained applications.
@@
To overcome this limitation of diffusion models, we adapt rectified flows [17]: a deterministic ODE-based analogue of diffusion that “straightens” transport paths in probability space. Rather than small random increments, rectified flows solve an ODE whose velocity field is trained to match the diffusion model’s instantaneous score, but whose trajectories remain nearly linear. This structure permits large time integration, reducing the number of solver steps by an order of magnitude or more, while provably recovering the same target distribution.
Hence, the main contributions of our paper are,
  • Rectified flow framework: We propose a deterministic transport approach for conditional PDE data, leveraging “straightened” trajectories that can be integrated in far fewer steps than typical diffusion-based methods. This reduced order model perspective challenges current wide-spread practices which rely on overtly-complicated interpolations between base and target measures.
  • Empirical gains in multi-scale problems: Across diverse 2D benchmarks for the Incompressible Navier–Stokes and Compressible Euler equations governing fluid dynamics, our method generates samples up to 22x faster than diffusion baselines (e.g. GenCFD), while retaining multi-scale accuracy. Furthermore, we observe a significant gain of efficiency in training, where rectified-flows converge reliably faster across all experimental testbeds, while preserving their strong performance.
@@ 2 Approach
[…]
subject to boundary conditions $\mathcal{B}(u)=u_{b}$ on $\partial D\times(0,T)$ and initial data $u(x,0)=\bar{u}(x)$. Here, the solution $u:D\mapsto{\mathbb{R}}^{m}$ is evolved in time, with respect to the differential operator $\mathcal{L}$. Concrete examples of (2.1) are the incompressible Navier-Stokes and Compressible Euler equations.
@@
Rectified flows for faster sampling. To mitigate these inefficiencies, we adopt the framework of rectified flows [17]: a deterministic alternative to conditional diffusion designed to reduce the number of required integration steps. In essence, rather than generating samples via small random “diffusion-reversal” increments, rectified flows construct an ordinary differential equation (ODE) with nearly straight trajectories in function space. This straighter transport map can be integrated more aggressively in time, enabling sampling with far fewer solver steps. Importantly, rectified flows retain the desirable property of matching the same final distributions that standard diffusion models learn, but at a fraction of the runtime cost.
Figure 2.1: Overview of ReFlow. (a): ReFlow closely matches high-order statistical moments (in the Wasserstein $W_{1}$-sense) of fluid flows by learning a neural surrogate to approximate the push-forward of the distribution $p(u_{\tau|\tau=0}|u_{i})$, while minimizing inference-time computational costs. (b) Macro-architecture overview: A UViT based network predicts the velocity field $v_{\theta}$, conditioned on Gaussian noise, initial data $u_{i}$ and boundary conditions, and diffusion time $\tau$. Noise and initial / boundary conditions are concatenated as inputs. (c): Results for the cylindrical shear flow dataset. (d): The KDE plots monitor the evolution of a down-projected version of the push-forward measure $p_{\tau}:=X_{\theta}(\tau)_{\#}p(u_{\tau=0}|u_{i})$, where $X_{\theta}$ is the flow induced by the learned vector field $v_{\theta}$. Results for the Richtmyer Meshkov dataset showing flow matching between noise ($p_{\tau=0}$) and target ($p_{\tau=1}$) distributions. ReFlow solves an ODE, learning a constant velocity field that enables straight-line inference paths (visualized with principal component analysis and kernel density estimation at different diffusion times).
@@ 3.1 Diffusion Models in Fluid Dynamics
Conditional diffusion surrogates. GenCFD [21] trains a conditional score-based diffusion model that starts from an initial or low-resolution flow field and iteratively refines it through hundreds of denoising steps, thereby capturing multi-scale turbulent structure. Similarly, the method of [23] conditions a diffusion model on a coarse fluid solution (generated by a learned operator) and “super-resolves” the fine-scale flow. Given these similarities, we refer to such diffusion-based solvers collectively as conditional diffusion PDE solvers.
Rectified flow alternative. Each sample typically needs hundreds of network calls, limiting large-scale or real-time use. Our method replaces the stochastic SDE with a deterministic ODE whose nearly straight trajectories require far fewer steps, achieving similar statistical fidelity at much lower inference cost (see Sections 4.3 and 5). Crucially, we find that rectified flows are actually more efficient from the standpoint of training-time required to reach the same accuracy level with the conditional diffusion approaches; this is also confirmed by earlier work [3] exploring the scalability of rectified flows with transformer-based backbones.
@@ 4.1 Mathematical Formulation
Throughout, our random variables are velocity fields $u:\Omega\!\to\!\mathbb{R}^{3}$ living in a separable, divergence-free space $\mathcal{U}=L^{2}_{\!\mathrm{div}=0}(\Omega;\mathbb{R}^{3})$ (or $H^{s}_{\!\mathrm{div}=0}$, $s>\frac{3}{2}$). All theory below extends verbatim from $\mathbb{R}^{d}$ because $\mathcal{U}$ is Polish; in practice we work with a Galerkin discretisation $\mathbb{R}^{d_{N}}\subset\mathcal{U}$, so every integral is evaluated by Monte-Carlo on its finite grid surrogate.

### v2

@@ Abstract
Statistical surrogate modeling of fluid flows is challenging due to multiscale dynamics and strong sensitivity to initial conditions. Conditional diffusion models can achieve high fidelity, but typically require hundreds of stochastic steps at inference. We introduce a rectified-flow surrogate that learns a time-dependent conditional velocity field transporting input-to-output laws along nearly straight trajectories. Sampling reduces to solving a deterministic ODE along this learned transport, so each function evaluation is substantially more effective: on multi-scale 2D benchmarks we match diffusion-class posterior statistics with as few as $8$ ODE steps versus $\geq\!128$ steps for score-based diffusion.
On the theory side, we develop a law-level analysis tailored to conditional PDE forecasts. First, we formalize the link between our evaluation criterion—one-point Wasserstein distances on fields—and the k=1k\!=\!1 correlation-marginal viewpoint in statistical solutions. Second, we provide a one-step error decomposition for the learned pushforward law into a coverage (high-frequency tail) term controlled by structure functions (equivalently, by spectral decay), and a fit term controlled directly by the training objective. Third, we show how straightness in rectification time governs local truncation error for ODE sampling, yielding step-count requirements and explaining why rectified transports admit large, stable steps.
Guided by this picture, we propose a curvature-aware sampler that tracks an EMA-based straightness proxy and adaptively blends and steps the velocity during inference. Across multiscale incompressible and compressible 2D flows, our method matches diffusion models in Wasserstein statistics and energy spectra, preserves fine-scale structure missed by MSE baselines, and delivers high-resolution conditional samples at a fraction of the inference cost.
@@ 1 Introduction
Partial differential equations (PDEs) Evans (2010) are the backbone of modeling physical phenomena, from atmospheric dynamics to aerodynamics and magnetohydrodynamics. Currently, solutions of PDEs are simulated with high-fidelity numerical solvers such as finite difference, finite element, and spectral methods Quarteroni and Valli (1994). These are often prohibitively expensive for many–query tasks Quarteroni et al. (2015) such as uncertainty quantification, optimization, and real‐time prediction. In particular, these solvers encounter a significant challenge for classes of non-linear PDEs, whose solutions exhibit extreme sensitivity to initial and boundary data, leading to chaotic behavior and complex multi‐scale features. In such settings, one seeks, not a deterministic trajectory, but rather the statistical solution of the PDE, i.e., the push-forward of the underlying probability measure on inputs (initial and boundary conditions etc) by the PDE solution operator Fjordholm et al. (2017). Currently, these statistical solutions are sampled with ensemble Monte Carlo method by repeatedly calling high‐fidelity PDE solvers Lanthaler et al. (2021b). Given the computational cost of the PDE solver, ensemble methods become infeasible as ensemble sizes grow.
@@
Machine learning and AI based algorithms are increasingly being explored for the data-driven simulation of PDEs Mishra and Townsend (2024). In particular, neural operators Kovachki et al. (2023), which directly seek to learn the PDE solution operator from data, are already widely used in a PDE context Mishra and Townsend (2024). Examples of neural operators have been proposed in Li et al. (2020a; b); Lu et al. (2021); Wu et al. (2024); Hao et al. (2023); Li et al. (2023); Alkin et al. (2024) and references therein.
@@
However, these neural operators have been found to be unsuitable for simulating PDE with chaotic, multiscale solutions such as governing equations of fluid dynamics. For instance, recently in Molinaro et al. (2024), it was shown both theoretically and empirically that neural operators, trained to minimize mean absolute (square) errors, will always collapse to the ensemble mean when used for simulating statistics of PDEs with chaotic multiscale solutions.
@@
Given this context, generative AI models—especially score‐based and diffusion models—have recently shown remarkable success at learning complex, high‐dimensional distributions by progressive denoising score matching Song et al. (2020). In the context of chaotic PDEs, GenCFD Molinaro et al. (2024) introduced a conditional diffusion framework that preserves spectral content across fine scales, mitigating the “spectral collapse” of naive MSE‐trained surrogates, see also Oommen et al. (2024); Gao et al. (2024a; b; c). However, sampling a score-based diffusion model requires simulating a reverse‐time stochastic differential equation (SDE) over 100+ discretization steps, each invoking a large ML model on a moderate-to-high resolution spatial grid, resulting in high runtimes that undermine real‐time or resource‐constrained applications.
@@
To overcome this limitation of diffusion models, we adapt rectified flows Liu et al. (2022): a deterministic ODE‐based analogue of diffusion that “straightens” transport paths in probability space. Rather than small random increments, rectified flows solve an ODE whose velocity field is trained to match the diffusion model’s instantaneous score, but whose trajectories remain nearly linear. This structure permits large time integration, reducing the number of solver steps by an order of magnitude or more, while provably recovering the same target distribution.
What is evaluated, and why.
Because statistical solutions are measures on function spaces, the most robust comparison is via observables or correlation marginals Lanthaler et al. (2021a). Our evaluation metric in Section 6.2—the one-point Wasserstein distance—is exactly the Wasserstein discrepancy between k=1k\!=\!1 correlation marginals (pointwise push-forwards), averaged over space. This aligns the experimental metrics (Table 1) with the law-level theoretical statements in Sections 4.14.3.
Hence, our main contributions are:
  • Rectified flows for conditional PDEs. We learn a time-dependent velocity field that transports a conditional law of fields from a simple reference (Gaussian noise conditioned on the input) to the target PDE pushforward law. The key theoretical point is correctness at the level of laws: the barycentric rectified velocity induces an ODE whose flow maps the initial law to the target law (SM C).
  • Curvature-aware integration. The sampling cost is governed by the amount of time-bending (straightness/curvature in rectification time) of the learned vector field along typical trajectories. We formalize: (i) which curvature quantity controls discretization error, (ii) why a simple EMA deviation is a principled proxy for that curvature, and (iii) why the blending step is the minimizer of a quadratic (Tikhonov) correction (SM E).
  • Multi-scale accuracy at speed. At fixed grid resolution there is an irreducible law-level error floor coming from unresolved Fourier scales. We express this as a coverage term controlled by structure functions / spectral tails (Crefapp:sec:coverage-structure). Rectified flows reduce curvature so that the best attainable (bandlimited) law is reached with far fewer sampling steps. This is precisely what is seen in Table 1 and in the spectral evolution plots (Figure 13).
@@ 2 Approach
$\partial_{t}\,u(x,t)\;+\;\mathcal{L}\!\bigl(u,\nabla_{x}u,\nabla_{x}^{2}u,\dots\bigr)\;=\;0,\quad x\in D\subset\mathbb{R}^{d},\;t\in(0,T),$ (1)
subject to boundary conditions $\mathcal{B}(u)=u_{b}$ on $\partial D\times(0,T)$ and initial data $u(x,0)=\bar{u}(x)$. Here, the solution $u:D\mapsto{\mathbb{R}}^{m}$ is evolved in time, with respect to the differential operator $\mathcal{L}$. Concrete examples of (1) are the incompressible Navier-Stokes and compressible Euler equations.
@@
Observables and correlation marginals (what metrics actually probe). A classical way to formalize “quantities of interest” for $\mu_{t}$ is via observables and their induced correlation marginals Lanthaler et al. (2021a). For $k=1$, the pointwise evaluation map at $x\in D$,
$\mathrm{ev}_{x}:\;u\mapsto u(x)\in\mathbb{R}^{m},$
pushes $\mu_{t}$ to the one-point marginal $(\mathrm{ev}_{x})_{\#}\mu_{t}$ on $\mathbb{R}^{m}$ – strictly speaking, one studies the pushforwards measures $\mathrm{ev}_{x}\circ i_{\epsilon}$, here $(i_{\epsilon}u):=(\phi_{\epsilon}\ast u)$ where $\phi_{\epsilon}$ is a compactly supported mollifier, so that point-wise evaluation is well-defined; for the sake of brevity, we will not make this distinction explicit in the sequel. Our one-point Wasserstein metric (Section 6.2) is the spatial average of $W_{1}\big((\mathrm{ev}_{x})_{\#}\mu_{t},(\mathrm{ev}_{x})_{\#}\widehat{\mu}_{t}\big)$. By Kantorovich–Rubinstein duality, controlling this quantity controls the discrepancy of all pointwise 1-Lipschitz observables, averaged over space. This is the bridge between Table 1 and the law-level bounds in Section 4.1 and Section 4.3.
Rectified flows for faster sampling. To mitigate these inefficiencies, we adopt the framework of rectified flows Liu et al. (2022): a deterministic alternative to conditional diffusion designed to reduce the number of required integration steps. In essence, rather than generating samples via small random “diffusion-reversal” increments, rectified flows construct an ordinary differential equation (ODE) with nearly straight trajectories in function space. This straighter transport map can be integrated more aggressively in time, enabling sampling with far fewer solver steps. Importantly, rectified flows retain the desirable property of matching the same final distributions that standard diffusion models learn, but at a fraction of the runtime cost.
Figure 1: Overview of ReFlow. (a): ReFlow closely matches high-order statistical moments (in the Wasserstein $W_{1}$-sense) of fluid flows by learning a neural surrogate to approximate the push-forward of the distribution $p(u_{\tau|\tau=0}|u_{i})$, while minimizing inference-time computational costs. (b) Macro-architecture overview: A UViT based network predicts the velocity field $v_{\theta}$, conditioned on Gaussian noise, initial data $u_{i}$ and boundary conditions, and diffusion time $\tau$. Noise and initial / boundary conditions are concatenated as inputs. (c): Results for the cylindrical shear flow dataset. (d): The KDE plots monitor the evolution of a down-projected version of the push-forward measure $p_{\tau}:=X_{\theta}(\tau)_{\#}p(u_{\tau=0}|u_{i})$, where $X_{\theta}$ is the flow induced by the learned vector field $v_{\theta}$. Results for the Richtmyer Meshkov dataset showing flow matching between noise ($p_{\tau=0}$) and target ($p_{\tau=1}$) distributions. ReFlow solves an ODE, learning a constant velocity field that enables straight-line inference paths (visualized with principal component analysis and kernel density estimation at different diffusion times).
@@ 3.1 Diffusion Models in Fluid Dynamics
Conditional diffusion surrogates. GenCFD Molinaro et al. (2024) trains a conditional score-based diffusion model that starts from an initial or low-resolution flow field and iteratively refines it through hundreds of denoising steps, thereby capturing multi-scale turbulent structure. Similarly, the method of Oommen et al. (2024) conditions a diffusion model on a coarse fluid solution (generated by a learned operator) and ”super-resolves” the fine-scale flow. Given these similarities, we refer to such diffusion-based solvers collectively as conditional diffusion PDE solvers.
Rectified flow alternative. Each sample typically needs hundreds of network calls, limiting large-scale or real-time use. Our method replaces the stochastic SDE with a deterministic ODE whose nearly straight trajectories require far fewer steps, achieving similar statistical fidelity at much lower inference cost (see Sections 5.6 and 6). Crucially, we find that rectified flows are actually more efficient from the standpoint of training-time required to reach the same accuracy level with the conditional diffusion approaches; this is also confirmed by earlier work Esser et al. (2024) exploring the scalability of rectified flows with transformer-based backbones.
@@ 4.1 Mathematical Formulation
Throughout, our random variables are state fields $u:\mathcal{D}\!\to\!\mathbb{R}^{m}$ (e.g. $m=2$ for velocity, or $m=4$ for compressible state variables) living in a separable space $\mathcal{U}=L^{2}(\mathcal{D};\mathbb{R}^{m})$ (or $H^{s}$, $s>\frac{d}{2}$). All theory below extends verbatim from $\mathbb{R}^{d}$ because $\mathcal{U}$ is Polish; in practice we work with a Galerkin discretisation $\mathbb{R}^{d_{N}}\subset\mathcal{U}$, so every integral is evaluated by Monte-Carlo on its finite grid surrogate.